CON705-32209-SPRING2019 e/My Courses/ECON705-32209-SPRING2019Module 1- Introduction, Supply & Demand / Module One: Assessment Qu The following represents demand for widgets (a fictional product) OD -700-100P+0.5M +30P where P is the price of widgets, M is income, and PR is the price of a related (fictional) good, the wodget. Supply of widgets is determined by Qs 900+57.5P Widgets are out of and widgets and wodgets are Select one: a. a normal good; substitutes. b. an inferior good; substitutes. c. a normal good; complements. d. an inferior good; complements. Assume that M $61, 000 and PR $250. The equilibrium price of widgets is 9 Answer t of
Answer #1

Correct Answer:

A

Explanation:

As per the given demand equation:

Increase in income, will lead to increase in quantity demanded and vice versa. So, the product is normal good. Further, increase in price of the related good (wodgate) leads to increase in demand of widget. So, widgets and wodgets are substitute to each other.

----------------------

Correct Answer:

Qd = 700-100P + .5*61000 + 30*250

Qd = 38700 - 100P

Qs = 900 + 57.5P

At equilibrium,

38700 - 100P = 900 + 57.5P

P = (38700-900)/(100+57.5)

P = 240

So,

Q = 38700 - 100*240

Q = 14700